Company A owns 40% of Company B
​Company B owns 30% of Company C
​Company C owns 20% of Company A
​Each company has a total market capitalization of $200 million. You wish to adjust for cross-holding to reflect the weights of these companies in a market weighted index. What adjustment would you make to reflect the free float?
Answer #1
Company A B C
20% owned by C 40% owned by A 30% owned by B
So, adjustment to reflect free float-- 200*(1-20%)= 200*(1-40%)= 200*(1-30%)=
out of total market-cap of $ 200 mlns. 160 120 140
so,total adjusted market capitalisation is 160+120+140= 420 millions
or alternatively,
Total mkt. capitalisation-Total crossholdings
(200*3)-(200*(40%+30%+20%))=
420
millions